solve by the method of your choice.
-x + y =1
2x - 3y = -7

Found 2 solutions by ewatrrr, mananth:
Answer by ewatrrr(24785)


Hi
Solving by substitution
-x + y =1 OR y = x+1
2x - 3y = -7
2x - 3(x+1) = -7
-x -3 = -7
x = 4 and y = 5 (4+1)


Answer by mananth(16949)

-x + y =1............................1
2x - 3y = -7..........................2
...
multiply (1) by 2
-2x+2y=2
2x-3y=-7
add
-y=-5
y=5
...
-x+y =1
y =5
so x=4
...